OpenStudy (anonymous):

-6.(x-4)-17=4(x+2)+19=

OpenStudy (debbieg):

are you solving -6.(x-4)-17=4(x+2)+19? If so, just multiply out the factors and then add/subtract terms to isolate variable, and solve: -6x+24-17=4x+8+19 -6x+7=4x+27 -10x=20 x=-2